OpenStudy (anonymous):

Solve for m in , if D= 5.1 and v = 0.3.

OpenStudy (.cheesecake.):

Hi :) Is there an equation that goes with the problem?

OpenStudy (anonymous):

nope

OpenStudy (.cheesecake.):

Solve for m in _____? in what?

OpenStudy (calculusxy):

I am pretty sure that after "in" there is supposed to be an equation.

OpenStudy (anonymous):

d = m/v

OpenStudy (calculusxy):

oh the density formula

OpenStudy (anonymous):

multiply both sides by v

OpenStudy (anonymous):

m=dv. now substitute values and solve

OpenStudy (anonymous):

m=(5.1)*(0.3)=?

OpenStudy (bubblegum.):

hey oleg come to the chess match! :D

OpenStudy (anonymous):

buble, lol, i forgot about you, sure xD

OpenStudy (.cheesecake.):

@OlegPetrovich_PR It would help if you knew how to type in one comment.

OpenStudy (koikkara):

then, D = m / v D = 5.1 v = 0.3 5.1 = m / .3 5.1 * .3 = m .153 = m

OpenStudy (koikkara):

@akroayee Done, now make sure you post the question good.

OpenStudy (anonymous):

yes